### v2

# Transversely projective structures on smooth foliations on surfaces

Abstract.
Brunella’s classification implies that every smooth foliation on a compact complex surface admits a singular transversely projective structure. However, Biswas and Dumitrescu’s recent work shows that certain foliations on compact complex surfaces, despite possessing a singular transversely projective structure, do not admit a regular transversely projective structure. Here, we describe the smooth foliations on compact complex surfaces which fail to possess a regular transversely projective structure.
Key words and phrases:
Holomorphic foliation; transversely projective structure; fibration
2020 Mathematics Subject Classification:
32M25, 37F75, 53C12
1. Introduction
In [3], Biswas and Dumitrescu employed a dimension counting argument to establish that a sufficiently general smooth foliation on the product of elliptic curves does not admit a transversely projective structure. Inspired by their work, our pourpose here is to describe all smooth holomorphic foliations on compact complex surfaces which do not admit a transversely projective structure.
Naturally, we build on Brunella’s classification of smooth foliations on compact complex surfaces. According to [4], if $X$ is a compact Kähler surface and $\mathcal{F}$ is a smooth foliation on it then $\mathcal{F}$ is a fibration, or $\mathcal{F}$ is transverse to the fibers of a $\mathbb{P}^{1}$-bundle, i.e. $\mathcal{F}$ is a Riccati foliation, or $\mathcal{F}$ is transverse to the general fiber of an elliptic fibration, i.e. $\mathcal{F}$ is a turbulent foliation, or $\mathcal{F}$ is a linear foliation on a complex torus, or $\mathcal{F}$ is a transversely hyperbolic foliation with dense leaves. Moreover, if $X$ is a non-Kähler compact surface carrying a smooth foliation $\mathcal{F}$ then $X$ is a Hopf surface or an Inoue surface and $\mathcal{F}$ is any foliation on them.
A careful look at Brunella’s list just present reveals that besides fibrations, turbulent foliation, and foliations on Hopf surfaces, every other smooth foliation on a compact complex surface admit a natural transversely projective structure.
As already pointed out by Biswas and Dumitrescu, turbulent foliations are not necessarily transversely projective. In Section 4 below, we characterize which turbulent foliations admit transverse projective structures, see for instance Theorem 4.4.
Curiously, there are fibrations on Hopf surfaces which do not admit a transversely projective structure. We show that this phenomenon does not happen when the surface is Kähler and compact, see Theorem 3.4. We also obtain an analogue statement for fibrations on projective manifolds of dimension three, see Theorem 3.8 for a precise statement. Although the general foliation on a Hopf surface is not turbulent, as it turns out, their analysis is very similar to the analysis of turbulent foliations as explained in Section 5.
Acknowledgements
We are indebt to Sorin Dumitrescu for bringing the problem addressed in this article to our attention and for correspondence regarding it. Furthermore, the authors express their appreciation for the financial support provided by: CAPES-COFECUB program (project number: Ma 1017/24, Géométrie des équations différentielles et des variétés algébriques), funded by the French Ministry for Europe and Foreign Affairs, the French Ministry for Higher Education and Research, and CAPES; and CNPq Projeto Universal 408687/2023-1 "Geometria das Equações Diferenciais Algébricas". Pereira acknowledges the support from CNPq (Grant number 301683/2019-0), and FAPERJ (Grant number E26/200.550/2023). Fazoli acknowledges the support of FAPERJ (Grant number E26/201.353/2023). Melo acknowledges the support of CAPES-PROEX.
2. Transversely projective foliations
2.1. Foliations
A smooth holomorphic foliation $\mathcal{F}$ on a complex manifold $X$ is determined by an involutive locally free subsheaf $T_{\mathcal{F}}$ of $T_{X}$ (the tangent sheaf/bundle of $\mathcal{F}$) with cokernel $T_{X}/T_{\mathcal{F}}$ also locally free. The cokernel $T_{X}/T_{\mathcal{F}}$ will be denoted by $N_{\mathcal{F}}$ and is the normal bundle of $\mathcal{F}$. The rank of $T_{\mathcal{F}}$ is the dimension of the foliation, while the rank of $N_{\mathcal{F}}$ is the codimension of the foliation. The dual of $N_{\mathcal{F}}$ is the conormal bundle of $\mathcal{F}$ and will be denoted by $N^{*}_{\mathcal{F}}$, and the dual of $T_{\mathcal{F}}$ is the cotangent bundle of $\mathcal{F}$ and will be denoted by $\Omega^{1}_{\mathcal{F}}$.
2.2. Transversely projective structures
We recall one of the many equivalent definitions of transversely projective foliations, see for instance [16].
Definition 2.1.
A codimension one foliation $\mathcal{F}$ on a complex manifold $X$ is a (regular) transversely projective foliation if there exists
  1. (1)
    a $\mathbb{P}^{1}$-bundle $\pi:P\rightarrow X$ over $X$;
  2. (2)
    a codimension one smooth foliation $\mathcal{G}$ on $P$ everywhere transverse to the fibers of $\pi$; and
  3. (3)
    a holomorphic section $\sigma:X\rightarrow P$ transverse to $\mathcal{G}$ at every point, such that $\mathcal{F}=\sigma^{*}\mathcal{G}$.
The datum $\mathcal{P}=(\pi,\mathcal{G},\sigma)$ is called a (smooth) transversely projective structure of $\mathcal{F}$.
Two transversely projective structures $\mathcal{P}=(\pi:P\to X,\mathcal{G},\sigma)$ and $\mathcal{P}^{\prime}=(\pi^{\prime}:P^{\prime}\to X,\mathcal{G}^{\prime},\sigma^{\prime})$ for the same foliation $\mathcal{F}$ are equivalent if there exists an invertible morphism $\Phi:P\to P^{\prime}$ such that $\pi^{\prime}\circ\Phi=\mathrm{Id}$, $\Phi^{*}\mathcal{G}^{\prime}=\mathcal{G}$ and $\Phi^{*}\sigma^{\prime}=\sigma$.
If $X$ is a curve and $\mathcal{F}$ is the foliation with $T_{\mathcal{F}}=0$ and $N_{\mathcal{F}}=T_{X}$ then a transversely projective structure for $\mathcal{F}$ is nothing but a projective structure for the Riemann surface $X$ as explained in [15, Section 1.5].
The same foliation may admit several non-equivalent projective structures.
Example 2.2 (Linear foliations on complex tori).
If $X$ be a compact complex torus and $\mathcal{F}$ be a codimension one foliation defined by a non-zero section $\omega\in H^{0}(X,\Omega^{1}_{X})$. For any $\lambda\in\mathbb{C}$, we may consider the foliation $\mathcal{G}_{\lambda}$ on the trivial $\mathbb{P}^{1}$-bundle $P=X\times\mathbb{P}^{1}$ defined by the meromorphic $1$-form $d\log z-\omega+z\lambda\omega$, where $z$ is a projective coordinate on $\mathbb{P}^{1}$. If $\sigma:X\to P$ is the section $\sigma(x)=(x,1)$ then $\mathcal{F}=\sigma^{*}\mathcal{G}_{\lambda}$. In this way, we obtain a family of transversely projective structures $(P,\mathcal{G}_{\lambda},\sigma)$ for $\mathcal{F}$ parameterized by $\lambda\in\mathbb{C}$. Moreover, if $\lambda\neq\lambda^{\prime}$ then the corresponding transversely projective structures are non-equivalent.
The set of projective structures on a Riemann surface $X$ is parameterized by the affine space $H^{0}(X,(\Omega^{1}_{X})^{\otimes 2})$, see [15]. We recall below an analogue for transversely projective foliations of this classical fact.
Proposition 2.3.
Let $\mathcal{F}$ be a smooth codimension one foliation on a complex manifold $X$. The equivalence classes of transversely projective structures of $\mathcal{F}$ is parameterized by the kernel of morphism
$H^{0}(X,(N^{*}_{\mathcal{F}})^{\otimes 2})\to H^{0}(X,(N^{*}_{\mathcal{F}})^{\otimes 2}\otimes\Omega^{1}_{\mathcal{F}})$
induced by Bott’s partial connection. In particular, if $h^{0}(X,(N^{*}_{\mathcal{F}})^{\otimes 2})=0$ then $\mathcal{F}$ admits at most one transversely projective structure.
We refer to [3] for a proof of Proposition 2.3.
2.3. Distinguished local first integrals
Alternatively, a transversely projective structure for a foliation $\mathcal{F}$ can be defined in terms of distinguished local first integrals.
Definition 2.4.
Let $X$ be a complex manifold. Let $\mathcal{F}$ be a smooth foliation on $X$. A transversely projective structure for $\mathcal{F}$ is a collection $\{f_{i}:U_{i}\rightarrow V_{i}\subset\mathbb{P}^{1}\}_{i\in I}$ of holomorphic submersions $f_{i}$ such that
  1. (1)
    the collection of domains $U_{i}$ forms an open cover of $X$;
  2. (2)
    for every $i\in I$, the submersion $f_{i}$ define ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{U_{i}}}$, that is
    ${\left.\kern-1.2ptT_{\mathcal{F}}\vphantom{\big{|}}\right|_{U_{i}}}=\ker df_{i}:{\left.\kern-1.2ptT_{X}\vphantom{\big{|}}\right|_{U_{i}}}\to f_{i}^{*}T_{\mathbb{P}^{1}};\text{and}$
  3. (3)
    whenever $U_{i}\cap U_{j}\neq\emptyset$, there exists $f_{ij}\in\mathrm{Aut}(\mathbb{P}^{1})$ such that
    ${\left.\kern-1.2ptf_{i}\vphantom{\big{|}}\right|_{U_{i}\cap U_{j}}}=f_{ij}\circ{\left.\kern-1.2ptf_{j}\vphantom{\big{|}}\right|_{U_{i}\cap U_{j}}}\,.$
Definitions 2.1 and 2.4 are equivalent. To pass from the first to the second, one starts with a sufficiently fine open covering $\{U_{i}\}$ of $X$, consider $\mathbb{P}^{1}$-valued submersions $g_{i}:\pi^{-1}(U_{i})\to\mathbb{P}^{1}$ which are first integrals for ${\left.\kern-1.2pt\mathcal{G}\vphantom{\big{|}}\right|_{\pi^{-1}(U_{i})}}$ (they exist and any two differ by left composition by an element of $\operatorname{Aut}(\mathbb{P}^{1})$, because of the transversality between $\mathcal{G}$ and $\pi$), and consider the first integrals for ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{U_{i}}}$ defined by $g_{i}\circ\sigma$. Reciprocally, to pass from Definition 2.4 to Definition 2.1 one constructs the $\mathbb{P}^{1}$-bundle $P$ by gluing $U_{i}\times\mathbb{P}^{1}$ with $U_{j}\times\mathbb{P}^{1}$ using the transition function $f_{ij}$ and interpret the first integral $f_{i}$ as a section of the $\mathbb{P}^{1}$-bundle. For further details on the equivalence of the two definitions, the reader can consult [12].
2.4. Canonical projective structures on Riemann surfaces
Specializing Definition 2.4 to codimension one foliations on curves (that is, curves), one recovers the definition of a projective structure on a Riemann surface in terms of an atlas with transition functions taking values in $\operatorname{Aut}(\mathbb{P}^{1})$.
Example 2.5 (Canonical projective structure on a Riemann surface).
Every Riemann surface $X$ admits a projective structure induced by the uniformization theorem. Moreover, such projective structure is invariant by the automorphism group of $X$.
Proof.
Let $X$ be a Riemann surface and $\pi:\tilde{X}\rightarrow X$ its universal cover. According to the uniformization theorem, $\tilde{X}$ is either $\mathbb{P}^{1}$, $\mathbb{C}$ or $\mathbb{D}$. Let $\iota:\tilde{C}\hookrightarrow\mathbb{P}^{1}$ be the natural inclusion. By considering all local sections of $\pi$ and composing with $\iota$, we obtain a collection of embeddings $f_{i}:U_{i}\rightarrow\mathbb{P}^{1}$. The collection $\{f_{i}:U_{i}\to f_{i}(U_{i})\subset\mathbb{P}^{1}\}$ is a projective structure for $X$. The invariance by the automorphism group of $X$ is clear since $\operatorname{Aut}(\mathbb{D})$ and $\operatorname{Aut}(\mathbb{C})$ are subgroups of $\operatorname{Aut}(\mathbb{P}^{1})$. ∎
Following [15], we will call the projective structure given by Example 2.5, the canonical projective structure of $X$. Example 2.5 has two natural counterparts for codimension one foliations.
Example 2.6 (Smooth fibrations).
Let $B$ be a Riemann surface. Let $\mathcal{F}$ be a foliation defined by a smooth fibration $f:X\to B$, that is, every fiber of $f$ is smooth and none of them is multiple. Then, every projective structure for $B$ induces a smooth transversely projective structure for $\mathcal{F}$, given as the pullback of the projective atlas. Since, by Example 2.5, every Riemann surface admits a smooth projective structure, then $\mathcal{F}$ is transversely projective.
Example 2.7 (Foliations transverse to smooth fibrations).
Let $\mathcal{F}$ be a smooth codimension one foliation on a complex manifold $X$. Assume $X$ is endowed with a proper smooth fibration $\pi:X\to B$ with one-dimensional fibers such that $\mathcal{F}$ is everywhere transverse to the fibers of $\pi$. Then $\mathcal{F}$ is transversely projective.
Proof.
The transversality between $\mathcal{F}$ and $\pi$ impose strong restrictions on the foliation $\mathcal{F}$ and on the fibration $\pi$ as explained in [7, Chapter 5] which we summarize. Lifting paths on the basis $B$ along leaves of $\mathcal{F}$ one verifies that any two fibers of $\pi$ are isomorphic. Let us choose one of them, say $F_{b}=\pi^{-1}(b)$. Moreover, if $\tilde{B}$ is the universal covering of $B$ and $\tilde{X}$ is the fiber product $X\times_{B}\tilde{B}$ then $\tilde{X}$ is biholomorphic to the product $\tilde{B}\times F_{b}$ in such a way that $\tilde{\mathcal{F}}$, the pull-back of $\mathcal{F}$ to $\tilde{X}$, is the foliation defined the projection on the second the factor. The original manifold $X$ can be recovered as a quotient of $\tilde{X}$ by $\pi_{1}(B,b)$ acting on $\tilde{B}\times F$ through deck transformations on the first factor and through a representation $\rho:\pi_{1}(B,b)\to\operatorname{Aut}(F_{b})$.
The canonical projective structure on $F_{b}$ defines a projective structure on $\tilde{\mathcal{F}}$. Since the canonical projective structure on $F_{b}$ is invariant by automorphisms, the projective structure for $\tilde{\mathcal{F}}$ is invariant by the action of $\pi_{1}(B,b)$ and therefore descends to a projective structure on $\mathcal{F}$. ∎
2.5. Monodromy representation
Given a transversely projective structure $(\pi:P\to X,\mathcal{G},\sigma)$ for a foliation $\mathcal{F}$ on a complex manifold $X$ and a point $x\in X$, the lifting of paths on $X$ to the leaves of the foliation $\mathcal{G}$ defines an anti-representation
$\rho:\pi_{1}(X,x)\to\operatorname{Aut}(\mathbb{P}^{1})\,,$
which completely determines the isomorphism class of the $\mathbb{P}^{1}$-bundle $P$. It can also be obtained through the analytic continuation of the distinguished first integrals for the projective structure.
If one considers the étale covering $p:Y\to X$ of $X$ determined by $\ker\rho$, one obtains that $p^{*}\mathcal{F}$, the pull-back of $\mathcal{F}$ to $Y$, is defined by a submersion $f:Y\to\mathbb{P}^{1}$ which is covariant for the action of $\pi_{1}(X,x)/\ker\rho=\rho(\pi_{1}(X,x))$ on $Y$ that is
$f(\gamma\cdot y)=\rho(\gamma)(f(y))$
for every $\gamma\in\pi_{1}(X,x)/\ker\rho$.
2.6. Singular projective structures
Although our focus is on smooth transversely projective structures, it seems appropriate to recall the definition of singular transversely projective structures.
Definition 2.8.
A codimension one foliation $\mathcal{F}$ on a complex manifold $X$ is a singular transversely projective foliation if there exists
  1. (1)
    a $\mathbb{P}^{1}$-bundle $\pi:P\rightarrow X$ over $X$; and
  2. (2)
    a codimension one foliation $\mathcal{G}$ on $P$ transverse to the generic fiber of $\pi$; and
  3. (3)
    a meromorphic section $\sigma:X\dashrightarrow P$ generically transverse to $\mathcal{G}$
such that $\mathcal{F}=\sigma^{*}\mathcal{G}$. The datum $\mathcal{P}=(\pi,\mathcal{G},\sigma)$ is called a singular transversely projective structure of $\mathcal{F}$.
It is well-known, and easy to check using Brunella’s classification, that every smooth foliation on a compact complex surface admits a singular transversely projective structure. For more about singular transversely projective foliations/structures, we refer to [18] (the first work to consider the concept), [16] for a detailed discussion of the above definition, and [17] for a description of the structure of codimension one singular transversely projective foliations on projective manifolds of arbitrary dimension.
3. Fibrations
Let $X$ be a complex manifold and let $B$ be a smooth curve. We say that a proper fibration $f:X\to B$ is quasi-smooth if the support of every fiber of $f$ is smooth. In other words, for every $b\in B$, the divisor $f^{*}b\in\operatorname{Div}(X)$ is a positive multiple of a smooth and irreducible hypersurface $L_{b}$. Therefore we may write $f^{*}b=n_{b}L_{b}$ with $n_{b}\geq 1$. When $n_{b}>1$ then the fiber of $f$ over $b$ is a multiple fiber with multiplicity $n_{b}$.
3.1. Orbifold base
It is convenient to associate to any quasi-smooth fibration the $\mathbb{Q}$-divisor
$\Delta=\sum_{b\in B}\frac{n_{b}-1}{n_{b}}\cdot b\in\operatorname{Div}(B)\otimes\mathbb{Q}\,,$
and consider the pair $(B,\Delta)$ as the orbifold base of the fibration in the sense of [8, Section 1].
Notice that $f^{*}\Delta\in\operatorname{Div}(X)$ and satisfies
$f^{*}\Delta=\sum_{b\in B}(n_{b}-1)L_{b}\,.$
A standard computation, which traces back to Darboux, shows that $N^{*}_{\mathcal{F}}=f^{*}\omega_{B}\otimes\mathcal{O}_{X}(f^{*}\Delta)$ (see [5, Section 2.3, Example 5]).
An orbifold base $(B,\Delta)$ is called good (also called uniformizable) if there exists a curve $C$ and a discrete subgroup $G\subset\operatorname{Aut}(C)$ such that $B\simeq C/G$ and the ramification divisor of the quotient morphism $\pi:C\to C/G\simeq B$ equals $\pi^{*}\Delta$. We will call the triple $(C,G,\pi)$ an unfolding (also called uniformization) of $(B,\Delta)$. Otherwise, the orbifold base $(B,\Delta)$ is called bad.
Most orbifolds are good as established by Bundgaard-Nielsen-Fox theorem stated below. For a brief account of its history, the reader can consult [20, Theorem 1.4]. We also refer to [14] and [10, Chapter 2] for more information about orbifolds.
Theorem 3.1.
If $B$ is a compact Riemann surface and $(B,\Delta)$ is a bad orbifold then $B\simeq\mathbb{P}^{1}$. Moreover, $\Delta=\frac{n-1}{n}p$, for $n\geq 2$ and $p\in\mathbb{P}^{1}$, or $\Delta=\frac{n-1}{n}p+\frac{m-1}{m}q$ for distinct integers $n,m\geq 2$ and distinct points $p,q\in\mathbb{P}^{1}$.
3.2. Orbifold base of a transversely projective fibration is good
Proposition 3.2.
Let $f:X\to B$ be a quasi-smooth fibration and let $\mathcal{F}$ be the foliation defined by it. Then $\mathcal{F}$ admits a smooth transversely projective structure if, and only if, the orbifold base $(B,\Delta)$ is good.
Proof.
Assume first that the orbifold base of $f$ is good. Let $(C,G,\pi)$ be an unfolding of $(B,\Delta)$. The normalization of the fiber product $\tilde{X}=X\times_{B}C$ comes endowed with a smooth fibration $\tilde{f}:\tilde{X}\to C$. The canonical projective structure on $C$ induces a transversely projective structure on the foliation defined by $\tilde{f}$. Moreover, this transversely projective structure is invariant by the natural action of $G$ on $\tilde{X}$. Therefore, it descends to a smooth projective structure for $\mathcal{F}$.
Assume now that $\mathcal{F}$ admits a smooth transversely projective structure. Let $\rho:\pi_{1}(X)\to\operatorname{Aut}(\mathbb{P}^{1})$ be the monodromy representation of the structure. Selberg’s lemma implies the existence of a finite index normal subgroup $\Gamma\subset\pi_{1}(X)$ such that $\rho(\Gamma)$ is torsion free. Let $G=\pi_{1}(X)/\Gamma$ and consider the Galois covering $\pi:\tilde{X}\to X$ with group $G$ determined by $\Gamma$. Since the holonomy representation of leaves factor through the monodromy representation of the projective structure, the foliation $\pi^{*}\mathcal{F}$ is defined by a smooth fibration $\tilde{f}:\tilde{X}\to C$, for some curve $C$. Moreover, the group $G$ acts on $C$ in such a way that $C/G\simeq B$ and, $(C,G,C\to B)$ is an unfolding of $(B,\Delta)$. It follows that the orbifold base is good. ∎
Example 3.3 (Quasi-smooth fibrations which are not transversely projective foliations).
Let $\lambda\in\mathbb{C}^{*}$ be a complex number with modulus strictly smaller than one. Let $\mathbb{Z}$ act on $\tilde{X}=\mathbb{C}^{2}-\{(0,0)\}$ through $\gamma(x,y)=(\lambda^{n}x,\lambda^{m}y)$ where $n,m$ are distinct positive integers. Let $X=\tilde{X}/\mathbb{Z}$ be the associated Hopf surface. If $\gcd(n,m)=1$ then the meromorphic $f(x,y)=(x^{m}:y^{n})$ from $\tilde{X}$ to $\mathbb{P}^{1}$ defines a quasi-smooth fibration on $X$ with orbifold base
$\left(\mathbb{P}^{1},\frac{m-1}{m}p+\frac{n-1}{n}q\right)$
where $p=(0:1)$ and $q=(1:0)$. Since this orbifold base is bad, the foliation on $X$ is not transversely projective.
3.3. Quasi-smooth fibrations on compact Kähler surfaces
Theorem 3.4.
Let $X$ be a compact Kähler surface. If $f:X\to B$ is a quasi-smooth fibration then the orbifold base of $f$ is good.
Combining Proposition 3.2 with Theorem 3.4 we obtain the following statement which we record for later use.
Corollary 3.5.
Let $X$ be a compact Kähler surface. If $f:X\to B$ is a quasi-smooth fibration then, after passing to an étale covering of $X$, the fibration $f$ becomes a smooth fibration.
The proof of Theorem 3.4 relies on a topological observation which holds for compact complex manifolds of arbitrary dimension.
Lemma 3.6.
Let $X$ be a compact complex manifold. If $f:X\to B$ is a quasi-smooth fibration with bad orbifold base then the topological Euler characteristic of the fibers and of $X$ is zero.
Proof.
Theorem 3.1 implies that $B$ is equal to $\mathbb{P}^{1}$ and $\Delta=\frac{n-1}{n}p$ or $\Delta=\frac{n-1}{n}p+\frac{m-1}{m}q$ with $n\neq m$. Moreover, in the second case, after replacing $X$ by an étale covering of order $\gcd(n,m)$, we may assume that $\gcd(n,m)=1$. As before, let us denote by $L_{p}$, $L_{q}$ the supports of the multiple fibers. The open subset of regular values of $f$ will be denoted by $B^{\circ}$ and its pre-image under $f$ by $X^{\circ}$.
The support of a multiple fiber, say $L_{p}$, seen as leaf of the induced foliation, has finite holonomy and, by Reeb Local Stability Theorem, the smooth fibers of $f$ (which are all diffeomorphic to a general fiber $L$) are (topological) $n$-sheeted coverings of $L_{p}$. Therefore,
(3.1) $\chi_{top}(L)=n\chi_{top}(L_{p})=m\chi_{top}(L_{q}).$
Consider the local systems over $B^{0}$ given by $R^{\ell}f_{*}\mathbb{C}$. If $B^{\circ}\simeq\mathbb{C}$ then the simply-connectedness of $\mathbb{C}$ implies that these local systems have trivial monodromy. The same holds true when $B^{\circ}\simeq\mathbb{C}^{*}$ but now because the order of the monodromy around $p$ divides $n$ and around $q$ divides $m$.
Let $g:L\rightarrow L$ be a non-trivial deck transformation. By Lefschetz trace formula (see for instance [13, Theorem 2C.3]),
$\sum_{i=0}^{2\dim L}(-1)^{i}\mathrm{tr}\Big{(}g_{*}:H_{i}(L,\mathbb{C})\rightarrow H_{i}(L,\mathbb{C})\Big{)}=0.$
Observe that $g_{*}$ corresponds to the monodromy representation of a path $\gamma$ around $p$. Therefore, since the monodromy of the local systems $R^{\ell}f_{*}\mathbb{C}$ is trivial, $g_{*}=\mathrm{Id}$. Hence,
$\sum_{i=0}^{2\dim L}(-1)^{i}\mathrm{tr}\Big{(}g_{*}:H_{i}(L,\mathbb{C})\rightarrow H_{i}(L,\mathbb{C})\Big{)}=\sum_{i=0}^{2\dim L}(-1)^{i}h_{i}(L,\mathbb{C})=\chi_{top}(L)$
Hence, $\chi_{top}(L)=0$. Also by 3.1, $\chi_{top}(L_{p})=\chi_{top}(L_{q})=0$. This shows that every fiber has topological Euler characteristic equal to zero, as claimed.
To verify the vanishing of the topological Euler characteristic of $X$, first note that the additivity of the topological Euler characteristic implies
$\chi_{top}(X)=\chi_{top}(X^{\circ})+\chi_{top}(L_{p})+\chi_{top}(L_{q})\,.$
The vanishing of $\chi_{top}(X^{\circ})$ follows from the multiplicativity of the Euler characteristic for smooth fibrations. The vanishing of $\chi_{top}(X)$ follows. ∎
3.4. Proof of Theorem 3.4
Aiming at a contradiction, let us assume that the orbifold base of $f$ is bad. We will keep the notation of the proof of Lemma 3.6. Moreover, in the case $f$ has two multiple fibers we can assume, without loss of generalities, that the multiplicities of the fibers are relatively prime.
Let $\iota:L\rightarrow X$ be the inclusion of the general fiber $L$ of $f$. According to Deligne’s invariant subspace theorem, see for instance [21, Theorem 16.2], the image of $\iota^{*}:H^{1}(X,\mathbb{C})\rightarrow H^{1}(L,\mathbb{C})$ is the subspace invariant by the monodromy representation of the local system $R^{1}f_{*}\mathbb{C}$ defined on $B^{\circ}$. Because it is trivial (see proof of Lemma 3.6), we conclude that $\iota^{*}$ is surjective. Furthermore, $\iota^{*}$ defines a morphism of Hodge structures. Hence we get a surjective morphism
$\iota^{*}:H^{0}(X,\Omega^{1}_{X})\rightarrow H^{0}(L,\Omega^{1}_{L}).$
Because $L$ is of genus one, we have a nowhere vanishing $1$-form $\omega_{L}\in H^{0}(L,\Omega^{1}_{L})$.
Let $\omega\in H^{0}(X,\Omega^{1}_{X})$ such that $\iota^{*}\omega=\omega_{L}$ and consider the cohomology class of $\alpha=[\omega\wedge\overline{\omega}]\in H^{2}(X,\mathbb{C})$. Note that the contraction of $\alpha$ with the homology class of $L$ is equal to $\int_{L}\alpha$ and, therefore, non-zero. It follows that the same holds true for every other fiber of $f$ (multiple or not) since their homology classes all lie on the same line. Hence the $1$-form $\omega$ is non-zero in every fiber and it defines a foliation $\mathcal{G}$ completely transversal to $\mathcal{F}$. In particular, $T_{X}=T_{\mathcal{F}}\oplus T_{\mathcal{G}}$. Using [2, Theorem C], see also [6] and [9, Theorem 1.4], the universal cover of $\pi:\widetilde{X}\rightarrow X$ is the product of two simply-connected Riemann surfaces, and the foliation $\pi^{*}(\mathcal{F})$ is the fibration over one of them. This would imply that $B$ is a good orbifold, a contradiction. ∎
3.5. Quasi-smooth fibrations on surfaces with bad orbifold base
From the Theorem 3.4, we know that if $f:X\rightarrow\mathbb{P}^{1}$ is a quasi-smooth fibration with bad orbifold base, then $X$ must be a non Kähler compact surface. Moreover, in Example 3.3, we constructed examples of quasi-smooth fibrations on Hopf surfaces with bad orbifold base. From this, one can ask if there are other examples. The next proposition shows that this is the only case.
Proposition 3.7.
Let $X$ be a compact complex surface and $f:X\rightarrow\mathbb{P}^{1}$ be a quasi-smooth fibration with bad orbifold base. Then $X$ is a Hopf surface.
Proof.
Let $f:X\rightarrow\mathbb{P}^{1}$ be a fibration with a bad orbifold base $(\mathbb{P}^{1},\frac{m-1}{m}p+\frac{n-1}{n}q)$ (for the case $(\mathbb{P}^{1},\frac{m-1}{m}p)$ the computation is similar). By Lemma 3.6, $f$ is an elliptic fibration, and since the fibers are irreducible curves, they can not be $(-1)$-curves. That is, $f$ is a relatively minimal elliptic fibration.
The first step is to conclude that $\operatorname{kod}(X)=-\infty$. By [1, Chapter V, Proposition 12.5], $\operatorname{kod}(X)=-\infty$ if, and only if, $\delta(f)<0$, where $\delta(f)$ of a elliptic fibration $f:X\rightarrow B$ is defined as
$\delta(f):=\chi(X,\mathcal{O}_{X})+\left(2g(S)-2+\deg(\Delta)\right).$
In our situation, we have that $\chi(X,\mathcal{O}_{X})=deg(f_{*}\omega_{X|\mathbb{P}^{1}})$ by [1, Chapter III, Theorem 18.2], and by [1, Chapter III, Theorem 18.2], we have $deg(f_{*}\omega_{X|\mathbb{P}^{1}})=0$. Therefore,
$\delta(f)=\chi(X,\mathcal{O}_{X})-\frac{1}{m}-\frac{1}{n}=-\frac{1}{m}-\frac{1}{n}<0,$
and we conclude that $\operatorname{kod}(X)=-\infty$.
From Theorem 3.4 and Lemma 3.6, we see that $X$ must be a non Kähler surface. By [1, Chapter V, Theorem 18.6], it follows that $X$ is a Hopf surface.
3.6. Quasi-smooth fibrations on projective threefolds
It seems natural to wonder whether Theorem 3.4 holds for quasi-smooth fibrations over curves on arbitrary compact Kähler manifolds. We believe that this is the case, and we could check it for quasi-smooth fibrations on projective threefolds.
Theorem 3.8.
Let $X$ be a projective manifold of dimension three. If $f:X\to B$ is a quasi-smooth fibration then the orbifold base of $f$ is good.
Proof.
After applying the minimal model program (MMP) for codimension one smooth foliations on projective threefolds [19, Theorem 1.3] we can assume that every fiber of $f$ is minimal, i.e. does not contain a smooth curve isomorphic to $\mathbb{P}^{1}$ with self-intersection $-1$.
As in the proof of Theorem 3.4, we will assume that the orbifold base is bad and look for a contradiction.
Lemma 3.6 implies that every fiber of $f$ has zero topological Euler characteristic. An inspection of the Enriques-Kodaira classification [1, Chapter VI, Table 10] shows that the fibers of $f$ are
  1. (1)
    $\mathbb{P}^{1}$-bundles over elliptic curves when they have negative Kodaira dimension; or
  2. (2)
    bi-elliptic surfaces or abelian surfaces when they have Kodaira dimension zero; or
  3. (3)
    minimal properly elliptic surfaces when they have Kodaira dimension one.
If the fibers have negative Kodaira dimension then $X$ itself is a $\mathbb{P}^{1}$-bundle over a smooth projective surface $S$ and the fibration descends to a fibration on $S$. In this case, the result follows from Theorem 3.4.
If the fibers are abelian surfaces then we can argue as in the proof of Theorem 3.4, but now using $2$-forms instead of $1$-forms, to obtain a one-dimensional foliation $\mathcal{G}$ everywhere transverse to $\mathcal{F}$. It follows from [6] that the universal covering is the product $\mathbb{C}^{2}\times B=\mathbb{C}^{2}\times\mathbb{P}^{1}$ with $f$ lifting to the fibration defined by projection to $\mathbb{P}^{1}$. It follows that $f$ is good.
If the fibers are bi-elliptic surfaces or have Kodaira dimension one then there exists a smooth foliation $\mathcal{E}$ by elliptic curves tangent to the fibration. Since $\mathcal{E}$ is smooth, it must be an isotrivial fibration. Therefore we can apply [9, Theorem 1.4] to produce a finite étale covering $p:Y\to X$ such that $p^{*}\mathcal{E}$ is a locally trivial elliptic fibration (no multiple fibers). Again, the fibration $f$ descends to a fibration on a smooth projective surface and we can conclude using 3.4 as in the case of negative Kodaira dimension . ∎
Remark 3.9.
We only used the projectiveness of $X$ in order to apply the results on the MMP for codimension one foliations. It is very likely that the argument above can be adapted to show the result for compact Kähler threefolds.
4. Smooth Turbulent foliations on compact Kähler surfaces
In this section, we are going to study when a smooth turbulent foliation on a surface admits a transversely projective structure.
4.1. Definition
Let $X$ be a complex surface. A smooth foliation $\mathcal{F}$ on $X$ is called turbulent if there exists a proper fibration $f:X\rightarrow B$ such that the fibers of $f$ are curves of genus one, and $\mathcal{F}$ is completely transverse to the general fiber of $f$. We will say that $\mathcal{F}$ is turbulent with respect to $f$, and that $f$ is the reference fibration of $\mathcal{F}$.
It follows from the definition, that the reference fibration of a smooth turbulent foliation is automatically quasi-smooth. Indeed, as explained in [5, Section 4.3], if the support of a fiber of $f$ is not smooth then it is not irreducible and one of its components is invariant by $\mathcal{F}$. Moreover, since such component have negative self-intersection, Camacho-Sad formula implies that it must contain a singularity of the foliation. But this contradicts the smoothness of $\mathcal{F}$.
4.2. Turbulent foliations without invariant fibers
Proposition 4.1.
Let $\mathcal{F}$ be a turbulent foliation on a compact Kähler surface $X$ with reference fibration $f:X\to B$. If no fiber of $f$ is $\mathcal{F}$-invariant then $\mathcal{F}$ is transversely projective.
Proof.
If the fibration is smooth than the result follows from Example 2.7. If the fibration is only quasi-smooth then the natural orbifold structure on the base is good according to Theorem 3.4. Therefore, there exists an étale Galois covering $\pi:Y\to X$ induced by a ramified covering $C\to B$ such that the reference fibration of the turbulent foliation $\pi^{*}\mathcal{F}$ is smooth. Example 2.7 provides a transverse projective structure for $\pi^{*}\mathcal{F}$. Moreover, since such projective structure is obtained by extending the canonical projective structure on a fiber of $f$, it must be invariant by the action of the deck transformations of $\pi$. Hence it descends to a transversely projective structure for $\mathcal{F}$. ∎
4.3. Multiple and invariant fibers
Let $X$ be a smooth complex surface, let $\mathcal{F}$ be a smooth turbulent foliation on $X$, and let $f:X\to B$ be its reference fibration. Let $B^{\circ}\subset B$ be the set of regular values of $f$, and let $B^{\ast}\subset B^{\circ}$ be equal to $B^{\circ}$ minus the points corresponding to smooth $\mathcal{F}$-invariant fibers.
The transversality of $\mathcal{F}$ and the general fiber of $f$ implies that the restriction of $f$ to $\pi^{-1}(B^{\circ})$ is locally trivial. Moreover, over discs $\mathbb{D}\subset B^{\circ}$ we have that $f^{-1}(\mathbb{D})\simeq\mathbb{D}\times E$ and the foliation $\mathcal{F}$ is either conjugated to the horizontal foliation defined by $dz$, when none of the fibers are $\mathcal{F}$-invariant, or, there exists a meromorphic $1$-form $\alpha$ on $\mathbb{D}$ with polar divisor equal to the union of the points below the $\mathcal{F}$-invariant fibers such that ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{f^{-1}(\mathbb{D})}}$ is defined by $\alpha+dz$.
The description of $f$ and $\mathcal{F}$ around a multiple fiber of multiplicity $n$ is similar. If $\mathbb{D}\subset B$ is a disc with $\mathbb{D}^{*}\subset B^{\circ}$ then $f^{-1}(\mathbb{D})$ is the quotient of $\mathbb{D}\times E$ by the automorphism
$(x,z)\mapsto\left(\exp(2\pi\sqrt{-1}/n)x,z+\tau\right),$
where $\tau\in E$ is a torsion point of order $n$, according to [1, Section 13 of Chapter V]. The foliation $\mathcal{F}$ is like in the smooth case except for the extra condition, in the case $f^{-1}(0)$ is a $\mathcal{F}$-invariant fiber, that the $1$-form $\alpha$ is invariant by the autmorphism $x\mapsto\exp(2\pi\sqrt{-1}/n)x$.
Lemma 4.2.
Let $\mathcal{F}$ be a smooth turbulent foliation on a complex surface $X$. Then, after passing to a finite étale covering of $X$, the foliation is given by a closed meromorphic 1-form without zeroes.
Proof.
The discussion carried out above implies the existence of an open cover $\mathcal{U}=\{U_{i}\}$ of $B$ such that ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{f^{-1}(U_{i})}}$ is defined by a closed meromorphic $1$-form $\omega_{i}$. We can further assume that the integral of $\omega_{i}$ along generators of the homology group of the fibers does not vary with the base point. Moreover, we can also assume that this integral is equal to one at one of the generators. This rigidifies the choice of the $\omega_{i}$ in such a way that over non-empty intersections $U_{i}\cap U_{j}$, the $1$-forms $\omega_{i}$ and $\omega_{j}$ differ by multiplication of a locally constant function with values in {±$\{\pm 1\}$, or {±$\{\pm 1,\pm i\}$, or $\{\exp(2\pi\sqrt{-1}j/6;j=0\ldots 5\}$ depending on the automorphism group of the general fiber. In any case, we obtain a representation of $\pi_{1}(B)$ into a finite cyclic group. After doing a base change of $f:X\to B$ to the associated étale covering of $B$, it is clearly possible to define the transformed foliation by a closed meromorphic $1$-form. ∎
We will need the following fact concerning the non existence of first integral on neighborhoods of $\mathcal{F}$-invariant fibers of a turbulent foliation.
Lemma 4.3.
Let $\mathcal{F}$ be a turbulent foliation on a smooth complex surface $X$. Suppose that $\mathcal{F}$ leaves invariant some fiber of the reference fibration. Then $\mathcal{F}$ does not admit a non-constant meromorphic first integral.
Proof.
Let $f:X\to B$ be the reference fibration of $\mathcal{F}$ and let $g$ be a meromorphic first integral of $\mathcal{F}$. First observe that $g$ does not have indeterminacy points. Indeed, an indeterminacy point of $g$ would correspond a point in the intersection of several distinct leaves of $\mathcal{F}$. The smoothness of $\mathcal{F}$ prevents the existence of such point. Therefore, $g$ is not only a meromorphic map, but it is actually a morphism. Since it is constant along the leaves of $\mathcal{F}$, it must map the $\mathcal{F}$-invariant $F_{0}$ fiber to a point. Hence all the fibers sufficiently close to $F_{0}$ must also be mapped to a point. Since $\mathcal{F}$ is turbulent, this implies that $g$ is constant. ∎
4.4. Transversely projective structures
Theorem 4.4.
Let $\mathcal{F}$ be a smooth turbulent foliation on a compact Kähler surface $X$. Assume that $\mathcal{F}$ leaves invariant at least one fiber of the reference fibration. Then, $\mathcal{F}$ admits a transversely projective structure if, and only if, after a finite étale covering, $\mathcal{F}$ is given by a closed meromorphic 1-form $\omega$ of one of the following types
  1. (1)
    $\omega$ is without residues and $(\omega)_{\infty}=2\sum C_{i}$, where $C_{i}$ are irreducible smooth curves, $C_{i}\cap C_{j}=\emptyset,\forall i\neq j$; or
  2. (2)
    $\omega$ is logarithmic with residues of the form ±$\pm\lambda$, for some $\lambda\in\mathbb{C}^{*}$.
Proof.
Assume first that $\mathcal{F}$ is transversely projective. After passing to a finite étale cover we can assume that the reference fibration has no multiple fibers (Corollary 3.5) and that $\mathcal{F}$ is defined by a closed meromorphic $1$-form $\omega$ without zeros (Lemma 4.2). Let $C_{1},\ldots,C_{k}$ the $\mathcal{F}$-invariant curves that appear in $(\omega)_{\infty}$. Each $C_{i}$ corresponds to a fiber $f^{-1}(p_{i})$. Let $U\subset B$ be a simply connected open set of $B$, such that $p_{i}\in U$ for every $U$. Let $X_{U}=f^{-1}(U)$. By Ehresmann’s Theorem, $X_{U}$ is $C^{\infty}$-diffeomorphic to the product of an arbitrary fiber and $U$. In particular, $\pi_{1}(X_{U})$ is abelian. The foliation ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}$ is still turbulent and transversely projective, and we are going to prove that ${\left.\kern-1.2pt\omega\vphantom{\big{|}}\right|_{X_{U}}}$ is of type (1) or (2).
Let $\mathcal{P}=(\pi:Y\rightarrow X_{U},\mathcal{G},\alpha:X_{U}\rightarrow Y)$ be a projective structure for ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}$, and $\rho:\pi_{1}(X_{U})\rightarrow\mathrm{PSL}(2,\mathbb{C})$ be its monodromy representation. Because $\pi_{1}(X_{U})$ is abelian, the representation $\rho$, up to conjugation, fits into one the following descriptions.
  1. (i)
    The image of $\rho$ is contained in the subgroup $(\mathbb{C}^{*},\cdot)\subset\operatorname{Aut}(\mathbb{C})\subset\operatorname{Aut}(\mathbb{P}^{1})$. In this case, the foliation $\mathcal{G}$ is defined by a logarithmic 1-form with residues ±$\pm 1$ locally of the form $\frac{dz}{z}$, where $z$ is a coordinate on $\mathbb{P}^{1}$.
  2. (ii)
    The image of $\rho$ is contained $(\mathbb{C},+)\subset\operatorname{Aut}(\mathbb{C})\subset\operatorname{Aut}(\mathbb{P}^{1})$. In this case, $\mathcal{G}$ is given by a meromorphic 1-form without residues with is locally of the form $dz$.
  3. (iii)
    The image of $\rho$ has image contained in the subgroup of order four $\left\{\mathrm{id},z\mapsto-z,z\mapsto\frac{1}{z},z\mapsto-\frac{1}{z}\right\}\subset\operatorname{Aut}(\mathbb{P}^{1})$. In this case, $\mathcal{G}$ has admits a global meromorphic first integral locally of the form $f(z)=(z-\frac{1}{z})^{2}$.
In case (iii), ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}=\alpha^{*}(\mathcal{G})$ has meromorphic first integral contradicting Lemma 4.3. For the cases (i) and (ii), let $\eta$ be the meromorphic 1-form defining $\mathcal{G}$. Because $\alpha^{*}(\eta)$ defines ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}$, there exists a meromorphic function $F\in\mathbb{C}(X_{U})$ such that
$\alpha^{*}(\eta)=F\cdot{\left.\kern-1.2pt\omega\vphantom{\big{|}}\right|_{X_{U}}}.$
Because $\eta$ and $\omega$ are both closed, then $F$ is first integral for ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X_{U}}}$. Therefore, $F$ is constant by Lemma 4.3. The transversality between $\alpha$ and $\mathcal{G}$ implies that the order of poles and residues of ${\left.\kern-1.2pt\omega\vphantom{\big{|}}\right|_{X_{U}}}$ are the same as the ones of $\eta$. The result follows.
Suppose now that there exists a finite étale covering $\pi:\tilde{X}\to X$ such that ~$\tilde{\mathcal{F}}=\pi^{*}\mathcal{F}$ is defined by a closed meromorphic 1-form $\omega$ of type (1) or (2). Without loss of generality, we can assume that $\pi$ is Galois with Galois group $G$. In case (1), local primitives $\int\omega$ form a collection of distinguished first integrals for $\tilde{\mathcal{F}}$ showing that $\tilde{\mathcal{F}}$ is transversely projective. Similarly, in case (2), we consider the local first integrals defined by $\exp\left(\lambda^{-1}\int\omega\right)$ to produce a transversely projective structure for $\tilde{\mathcal{F}}$.
Since $\tilde{\mathcal{F}}$ is defined by a closed meromorphic $1$-form without zeros and with non-empty polar divisor, $N_{\tilde{\mathcal{F}}}=\mathcal{O}_{\tilde{X}}((\omega)_{\infty})$ is effective and non-zero. Therefore, $H^{0}((N^{*}_{\tilde{\mathcal{F}}})^{\otimes 2})=0$, and thus by Proposition 2.3, the transversely projective structure of $\tilde{\mathcal{F}}$ is unique. In particular, it is invariant by the Galois group $G$. Thus, the transversely projective structure descends to a transversely projective structure for $\mathcal{F}$. The theorem follows. ∎
The next example illustrates how the theorem above, applied in some special cases, provides examples of turbulent foliations that admit (and does not admit) transversely projective structures.
Example 4.5.
Let $E$ be an elliptic curve, and let $X=E\times\mathbb{P}^{1}$ with projections $\pi_{1}:X\rightarrow E$ and $\pi_{2}:X\rightarrow\mathbb{P}^{1}$. The smooth turbulent foliations on $X$ are given by closed meromorphic 1-forms of the form $\pi_{1}^{*}(\omega_{1})+\pi_{2}^{*}(\omega_{2})$, where $\omega_{1}$ is a holomorphic non-vanishing 1-form on $E$, and $\omega_{2}$ is a meromorphic 1-form on $\mathbb{P}^{1}$ (see [11, Theorem 1.2]). Then, $\mathcal{F}$ is smooth transversely projective if and only if $\omega_{2}$ satisfies one the conditions of Theorem 4.4. For instance, considering a coordinate $z$ for $\mathbb{P}^{1}$, both $\omega_{2}=dz$ and $\omega_{2}=dz/z$ provide examples of transversely projective foliations, while $\omega_{2}=dz/z^{3}$ provides an example of a turbulent foliation without transversely projective structure. More generally, one can verify that a generic smooth turbulent foliation in $X$ does not admit a transversely projective structure.
5. Foliations on Hopf and Inoue surfaces
5.1. Foliations on Inoue surfaces
An Inoue surface $X$ is a quotient of $\mathbb{H}\times\mathbb{C}$ by a (non-abelian) subgroup of $\operatorname{Aut}(\mathbb{H}\times\mathbb{C})\subset\operatorname{Aut}(\mathbb{C}^{2})$ generated by one diagonal linear map and three translations, see [1, Chapter V, Section 19]. The foliations defined by the natural projections to $\mathbb{H}$ and $\mathbb{C}$ descend to foliations on $X$ and are the unique foliations on $X$. Both are clearly transversely projective (actually, they are transversely affine).
5.2. Foliations on Hopf surfaces
A Hopf surface $X$ is a compact complex surface which has as universal covering $W=\mathbb{C}^{2}-\{(0,0)\}$, see [1, Chapter V, Section 18]. If $\pi_{1}(X)$ is isomorphic to $\mathbb{Z}$ then $X$ is called a primary Hopf surface, otherwise $X$ is a secondary Hopf surface. Any secondary Hopf surface admits a finite étale covering which is a primary Hopf surface. Primary Hopf surfaces in their turn, are $C^{\infty}$-diffeomorphic to $S^{3}\times S^{1}$ and biholomorphic to the cyclic quotient of $W$ by the group generated by either $\varphi(x,y)=(\alpha x,\beta y)$ where $\alpha,\beta$ are complex numbers satisfying $10<|\alpha|\leq|\beta|<$ 1 or $\varphi(x,y)\mapsto(\alpha^{n}x+\tau\cdot y^{n},\alpha y)$, where $\alpha$ is as before, $\tau\in\mathbb{C}^{*}$, and $n\geq 2$ is an integer.
5.3. Linear primary Hopf surfaces
Assume that we are in the first case, that is $\varphi:W\to W$ is a linear contraction. If the multiplicative semi-group of $\mathbb{C}^{*}$ generated by $\alpha$ and $\beta$ has rank two then the corresponding primary Hopf surface is of algebraic dimension zero and the foliations on it are quotients of the foliations defined by the $1$-forms $\lambda\frac{dx}{x}+\mu\frac{dy}{y}$ where $\lambda$ and $\mu$ are complex numbers, not both zero. If instead the multiplicative semi-group of $\mathbb{C}^{*}$ generated by $\alpha$ and $\beta$ has rank one, i.e. there exists relatively prime positive integers $n,m>0$ such that $\alpha^{n}=\beta^{m}$ then the algebraic dimension of $X$ is equal to one. Furthermore, the meromorphic function $x^{n}/y^{m}$ descends to a meromorphic function $f$ on $X$ and is such that $\mathbb{C}(X)=\mathbb{C}(f)$. Notice that $f$ has no indeterminacy points and defines a morphism from $X\to\mathbb{P}^{1}$ with elliptic fibers. Note that the orbifold structure on $\mathbb{P}^{1}$ is good only when $n=m=1$. In any case, every other foliation $\mathcal{F}$ on the primary Hopf surface $X$ is turbulent with respect to this fibration. Moreover, they are all defined by a closed meromorphic $1$-form $\omega$. We point out that $\mathcal{F}$ has at least one invariant fiber and hence, by Lemma 4.3, it does not have a meromorphic first integrals. Alternatively, this also follows from the fact the algebraic dimension of $X$ is equal to one. As a consequence, we obtain that the $1$-form $\omega$ is, up to multiplication by a non-zero constant, the unique closed meromorphic $1$-form defining $\mathcal{F}$.
5.4. Non-linear primary Hopf surfaces
We now discuss the second possibility for $\varphi$, that is $\varphi$ is non-linear and of the form $\varphi(x,y)\mapsto(\alpha^{n}x+\tau\cdot y^{n},\alpha y)$. In this case, the primary Hopf surface $X$ has algebraic dimension zero and the foliations on $X$ are obtained as quotients of the foliations on $W$ defined by the meromorphic $1$-forms
$\lambda\frac{dy}{y}+\mu d\left(\frac{x}{y^{n}}\right)\,,$
for $\lambda,\mu\in\mathbb{C}$. Notice that they are all closed and invariant by $\varphi$, thus they descend to closed meromorphic $1$-forms on $X$.
5.5. Synthesis
Having the description of the foliations on primary Hopf surfaces at hand, we can argue as in the proof of Theorem 4.4 to obtain the following result.
Proposition 5.1.
Let $\mathcal{F}$ be a foliation on a Hopf surface. Assume that the general leaf of $\mathcal{F}$ is not compact. Then $\mathcal{F}$ is transversely projective if, and only if, after a finite étale covering, $\mathcal{F}$ is given by a closed meromorphic 1-form $\omega$ of one of the following types
  1. (1)
    $\omega$ is without residues and $(\omega)_{\infty}=2\sum C_{i}$, where $C_{i}$ are irreducible smooth curves, $C_{i}\cap C_{j}=\emptyset,\forall i\neq j$; or
  2. (2)
    $\omega$ is logarithmic with residues of the form ±$\pm\lambda$, for some $\lambda\in\mathbb{C}^{*}$.
We provide an application of the preceding proposition, exhibiting cases of turbulent foliations that either possess or lack a transversely projective structure, depending on the choice of the defining 1-form.
Example 5.2.
Considering the description of the non-linear primary Hopf surfaces, Proposition 5.1 implies that a foliation $\mathcal{F}$ is transversely projective if, and only if, $\lambda=0$ and $n=1$.